Count the number of 6-digit numbers with:

  • Strictly ascending digits from left to right. For example 124579. Hint: any 6 distinct digits can be arranged in ascending order in a unique way.

- The answer is 84

  • Ascending digits from left to right. For example 133588, 566666, 788888. (hint: choose 6 digits out of {1,2,⋯,9} with repetition).

What ascending means is that each number is at least the previous.

- Please solve the second part of the question.

Ascending digits from left to right. For example 133588. (hint: choose 6 digits out of 1,2,...,9} with repetition)
Answer #1

Part b)

If repeated digits are allowed such that each digit is at least the previous one (eg 133588), is n+r-1Cr ­

So,

n = 9 and r = 6 (6 digit no)

n+r-1Cr ­ = 9+6-1C6 ­= 3003